# Bayesian Multiobject Tracking With Neural-Enhanced Motion and Measurement Models

### v2

@@ IV-B The Neural Architecture
“Encoder A” consists of a multilayer perceptron (MLP) with a gated recurrent unit (GRU). Based on the object state, $\bm{\mathbf{x}}^{i}_{k-1}$, and the hidden state, $\bm{\mathbf{h}}^{i}_{k-1}$, it produces a feature vector $f^{i}_{\bm{\mathbf{x}}_{k-1}}\in\mathbb{R}^{d_{\mathrm{h}}}$. Using the hidden state $\bm{\mathbf{h}}^{i}_{k-1}$ as an input to Encoder A is also an advancement over the RNN architecture that has led to significantly improved performance in our numerical evaluations. “Encoder B” is also a MLP. The state of neighbors indexed $m\in\{1,\dots,M\}$, i.e., each entry $s^{(i,m)}_{k-1}$, $m\in\{1,\dots,M\}$ in ${\mathbf{s}}^{i}_{k-1}$, is processed by this encoder, resulting in feature vectors $f^{(i,m)}_{\bm{\mathbf{s}}_{k-1}}\in\mathbb{R}^{d_{\mathrm{h}}}$.
@@
$f^{i}_{\bm{\mathbf{s}}_{k-1}}=\sum^{M}_{m=1}{w_{k}^{(i,m)}f^{(i,m)}_{\bm{\mathbf{s}}_{k-1}}}.\vspace{-.8mm}$ (27)
If there are fewer than $M$ neighbors, i.e., $M^{i}_{k}<M$, weights for $m>M^{i}_{k}$ are set equal to zero.
To compute an updated hidden state $\bm{\mathbf{h}}^{i}_{k}$, the feature vectors $f^{i}_{\bm{\mathbf{x}}_{k-1}}$ and $f^{i}_{\bm{\mathbf{s}}_{k-1}}$ are stacked and, together with the previous hidden state $\bm{\mathbf{h}}^{i}_{k-1}$, used as an input to the GRU, i.e.,
@@
$\bm{\mathbf{x}}^{i}_{k}=\mathrm{Decoder}(\bm{\mathbf{h}}^{i}_{k},\bm{\mathbf{x}}^{i}_{k-1}).\vspace{0mm}$ (28)
@@
Details on the loss function and training procedure are provided in the supplementary material [58].
@@ IV-C The Neural-Enhanced Prediction Step
The predicted PDF for each legacy PO is obtained via marginalization, i.e.,
@@
$\hskip 8.53581pt=\sum_{r^{i}_{k-1}\in\{0,1\}}\int\int\int f(\underline{\bm{\mathbf{x}}}^{i}_{k},\underline{r}^{i}_{k}|\bm{\mathbf{x}}^{i}_{k-1},r^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1})$
×$\hskip 8.53581pt\times f(\bm{\mathbf{y}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1}|\bm{\mathbf{z}}_{1:k-1})\mathrm{d}\bm{\mathbf{x}}^{i}_{k-1}\mathrm{d}{\mathbf{s}}^{i}_{k-1}\mathrm{d}\bm{\mathbf{h}}^{i}_{k-1}.$ (29)
For $\underline{r}^{i}_{k}\!=\!0$, we have $f(\underline{\bm{\mathbf{x}}}^{i}_{k},0\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1})=p\big{(}\underline{r}^{i}_{k}\!=0|\bm{\mathbf{z}}_{1:k-1}\big{)}$ $f_{\mathrm{d}}\big{(}\underline{\bm{\mathbf{x}}}^{i}_{k}\big{)}$, where $p\big{(}\underline{r}^{i}_{k}\!=0|\bm{\mathbf{z}}_{1:k-1}\big{)}$ can be obtained as
$p\big{(}\underline{r}^{i}_{k}\!=0|\bm{\mathbf{z}}_{1:k-1}\big{)}$
$\hskip 2.84526pt=\int f(\bm{\mathbf{x}}^{i}_{k-1},0|\bm{\mathbf{z}}_{1:k-1})+\!\big{(}1\!-p_{\mathrm{s}}\big{)}f(\bm{\mathbf{x}}^{i}_{k-1},1|\bm{\mathbf{z}}_{1:k-1})\,\mathrm{d}\bm{\mathbf{x}}^{i}_{k-1}$
$\hskip 2.84526pt=p\big{(}r^{i}_{k-1}\!=0|\bm{\mathbf{z}}_{1:k-1}\big{)}+\!\big{(}1\!-p_{\mathrm{s}}\big{)}p\big{(}r^{i}_{k-1}\!=1|\bm{\mathbf{z}}_{1:k-1}\big{)}.$ (30)
Here, we used that fact for $\underline{r}^{i}_{k}\!=0$, the state transition function in (29) simplifies as $f(\underline{\bm{\mathbf{x}}}^{i}_{k},\underline{r}^{i}_{k}\!=0|\bm{\mathbf{x}}^{i}_{k-1},r^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1})=f(\underline{\bm{\mathbf{x}}}^{i}_{k},\underline{r}^{i}_{k}\!=0|\bm{\mathbf{x}}^{i}_{k-1},r^{i}_{k-1})$ (cf. (7) and (8)).
For $\underline{r}^{i}_{k}\!=\!1$, the marginalization in (29) reads
[…]
$\hskip 2.84526pt=p\big{(}r^{i}_{k-1}\!=1|\bm{\mathbf{z}}_{1:k-1}\big{)}\hskip 0.85358pt\hskip 0.85358ptp_{\mathrm{s}}\int\int\int f({\underline{\bm{\mathbf{x}}}}^{i}_{k}|{\bm{\mathbf{x}}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1})$
×$\hskip 2.84526pt\times f({\bm{\mathbf{x}}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1},r^{i}_{k-1}\!=1)\hskip 0.85358pt\mathrm{d}\bm{\mathbf{x}}^{i}_{k-1}\mathrm{d}{\mathbf{s}}^{i}_{k-1}\mathrm{d}{\mathbf{h}}^{i}_{k-1}.$
@@
$f(\bm{\mathbf{h}}^{i}_{k}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1},r^{i}_{k-1}\!=1)$
@@
×$\hskip 1.42262pt\times f({\bm{\mathbf{x}}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1},r^{i}_{k-1}\!=1)\hskip 0.85358pt\mathrm{d}\bm{\mathbf{x}}^{i}_{k-1}\mathrm{d}{\mathbf{s}}^{i}_{k-1}\mathrm{d}{\mathbf{h}}^{i}_{k-1}.$
@@
To implement (LABEL:eq:prediction_1Prop) and (LABEL:eq:prediction_1Proph), we furthermore approximate the marginal posterior PDF $f({\bm{\mathbf{x}}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1})$ as
$f({\bm{\mathbf{x}}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1},r^{i}_{k-1}\!=1)$
[…]
×$\hskip 79.66771pt\times\prod^{M}_{m=1}f(\bm{\mathbf{s}}^{(i,m)}_{k-1}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1},r^{i}_{k-1}\!=1)$ (33)
$\approx\mathcal{N}\big{(}\bm{\mathbf{\mu}}_{{\bm{\mathbf{x}}}^{i}_{k-1}},\bm{\mathbf{C}}_{{\bm{\mathbf{x}}}^{i}_{k-1}}\big{)}\hskip 0.85358pt\delta\big{(}\bm{\mathbf{h}}^{i}_{k-1}-\hat{\bm{\mathbf{h}}}^{i}_{k-1}\big{)}$
×$\hskip 79.66771pt\times\prod^{M}_{m=1}\mathcal{N}\big{(}\bm{\mathbf{\mu}}_{{\bm{\mathbf{s}}}^{(i,m)}_{k-1}},\bm{\mathbf{C}}_{{\bm{\mathbf{s}}}^{(i,m)}_{k-1}}\big{)}.$ (34)
Computation of the point estimate $\hat{\bm{\mathbf{h}}}^{i}_{k-1}$ and the Gaussian parameters $\bm{\mathbf{\mu}}_{{\bm{\mathbf{x}}}^{i}_{k-1}}$, $\bm{\mathbf{C}}_{{\bm{\mathbf{x}}}^{i}_{k-1}}$, and $\bm{\mathbf{\mu}}_{{\bm{\mathbf{s}}}^{(i,m)}_{k-1}}$, $\bm{\mathbf{C}}_{{\bm{\mathbf{s}}}^{(i,m)}_{k-1}}$, $m\in\{1,\dots,M\}$ will be discussed in the last paragraph of this section. Based on the approximation in (34), we can implement (LABEL:eq:prediction_1Prop) and (LABEL:eq:prediction_1Proph) by means of SPs. An SP-based implementation is motivated by the nonlinear nature of the motion model $g(\bm{\mathbf{x}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1})$ and the fact that an implememtation based on particles would be computationally infeasible. Note that the fact that the joint PDF $f({\bm{\mathbf{x}}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1})$ fully factorizes as in (34), is not required for SP-based implement. This factorization naturally arises from the computation of marginal posterior PDFs by means of BP [4].
For SP-based implement of the high-dimensional integration in (LABEL:eq:prediction_1Prop) and (LABEL:eq:prediction_1Proph), we first define a stacked mean vector $\bm{\mathbf{\mu}}^{i}_{k-1}\in\mathbb{R}^{L(M+1)}$ and covariance matrix $\bm{\mathbf{C}}^{i}_{k-1}\in\mathbb{R}^{L(M+1)\times L(M+1)}$ corresponding to the product $\mathcal{N}\big{(}\bm{\mathbf{\mu}}_{{\bm{\mathbf{x}}}^{i}_{k-1}},\bm{\mathbf{C}}_{{\bm{\mathbf{x}}}^{i}_{k-1}}\big{)}\prod^{M}_{m=1}\mathcal{N}\big{(}\bm{\mathbf{\mu}}_{{\bm{\mathbf{s}}}^{(i,m)}_{k-1}},\bm{\mathbf{C}}_{{\bm{\mathbf{s}}}^{(i,m)}_{k-1}}\big{)}$ [59], i.e.,
@@
$\bm{\mathbf{C}}^{i}_{k-1}$ $\hskip 0.85358pt\triangleq\,\mathrm{bdiag}\Big{\{}\bm{\mathbf{C}}_{{\bm{\mathbf{x}}}^{i}_{k-1}},\bm{\mathbf{C}}_{{\bm{\mathbf{s}}}^{(i,1)}_{k-1}},\bm{\mathbf{C}}_{{\bm{\mathbf{s}}}^{(i,2)}_{k-1}},\ldots,\bm{\mathbf{C}}_{{\bm{\mathbf{s}}}^{(i,M)}_{k-1}}\Big{\}}.$
@@
From $\bm{\mathbf{\mu}}^{i}_{k-1}$ and $\bm{\mathbf{C}}^{i}_{k-1}$, $P=1+2L(M+1)$ weighted SPs $\big{\{}\big{(}\omega^{(i,p)}_{k},\bm{\mathbf{x}}^{(i,p)}_{k-1},{\mathbf{s}}^{(i,p)}_{k-1}\big{)}\big{\}}^{P}_{p=1}$ can be computed [12, 59]. SPs are next propagated through the augmented function, i.e, $\big{[}\bm{\mathbf{x}}^{(i,p)}_{k},{\mathbf{h}}^{(i,p)}_{k}\big{]}=g\big{(}\bm{\mathbf{x}}^{(i,p)}_{k-1},{\mathbf{s}}^{(i,p)}_{k-1},\hat{\bm{\mathbf{h}}}^{i}_{k-1}\big{)}$, $p\in\{1,\dots,P\}$. In this way, we obtain SPs $\big{\{}\big{(}\omega^{(i,p)}_{k},\bm{\mathbf{x}}^{(i,p)}_{k}\big{)}\big{\}}^{P}_{p=1}$ and $\big{\{}\big{(}\omega^{(i,p)}_{k},{\mathbf{h}}^{(i,p)}_{k}\big{)}\big{\}}^{P}_{p=1}$. Based on these SPs, a Gaussian approximation of $f(\underline{\bm{\mathbf{x}}}^{i}_{k},1|\bm{\mathbf{z}}_{1:k-1})$, is finally given by
$f(\underline{\bm{\mathbf{x}}}^{i}_{k},1|\bm{\mathbf{z}}_{1:k-1})\approx p\big{(}r^{i}_{k-1}\!=1|\bm{\mathbf{z}}_{1:k-1}\big{)}\hskip 0.85358ptp_{\mathrm{s}}\hskip 0.85358pt\hskip 0.85358pt\mathcal{N}\big{(}\bm{\mathbf{\mu}}_{{\underline{\bm{\mathbf{x}}}}^{i}_{k}},\bm{\mathbf{C}}_{\underline{{\bm{\mathbf{x}}}}^{i}_{k}}\big{)}.$
Here, mean and covariances computed as $\bm{\mathbf{\mu}}_{{\underline{\bm{\mathbf{x}}}}^{i}_{k}}=\sum_{p=1}^{P}\omega^{(i,p)}_{k}\bm{\mathbf{x}}^{(i,p)}_{k-1}\vspace{-.5mm}$ and $\bm{\mathbf{C}}_{\underline{{\bm{\mathbf{x}}}}^{i}_{k}}={\bm{\mathbf{Q}}}+\sum_{p=1}^{P}\omega^{(i,p)}_{k}(\bm{\mathbf{x}}^{(i,p)}_{k}\!-\bm{\mathbf{\mu}}_{{\underline{\bm{\mathbf{x}}}}^{i}_{k}})(\bm{\mathbf{x}}^{(i,p)}_{k}\!-\bm{\mathbf{\mu}}_{{\underline{\bm{\mathbf{x}}}}^{i}_{k}})^{\mathrm{T}}$, respectively. In addition, an updated point estimate $\hat{\bm{\mathbf{h}}}^{i}_{k}$ can be obtained as $\hat{\bm{\mathbf{h}}}^{i}_{k}=\sum^{P}_{p=1}\omega^{(i,p)}_{k}\hskip 0.85358pt\hskip 0.85358pt\bm{\mathbf{h}}^{(i,p)}_{k}.$

### v3

@@ IV-B The Neural Architecture
“Encoder A” consists of a multilayer perceptron (MLP) with a gated recurrent unit (GRU). Based on the object state, $\bm{\mathbf{x}}^{i}_{k-1}$, and the hidden state, $\bm{\mathbf{h}}^{i}_{k-1}$, it produces a feature vector $f^{i}_{\bm{\mathbf{x}}_{k-1}}\hskip-0.85358pt\in\hskip-0.85358pt\mathbb{R}^{d_{\mathrm{h}}}$. Using the hidden state $\bm{\mathbf{h}}^{i}_{k-1}$ as an input to Encoder A is also an advancement over the RNN architecture that has led to significantly improved performance in our numerical evaluations. “Encoder B” is also a MLP. The state of neighbors indexed $m\in\{1,\dots,M\}$, i.e., each entry $s^{(i,m)}_{k-1}$, $m\hskip-0.85358pt\in\hskip-0.85358pt\{1,\dots,M\}$ in ${\mathbf{s}}^{i}_{k-1}$, is processed by this encoder, resulting in feature vectors $f^{(i,m)}_{\bm{\mathbf{s}}_{k-1}}\hskip-0.85358pt\in\hskip-0.85358pt\mathbb{R}^{d_{\mathrm{h}}}$.
@@
[…]
If there are fewer than $M$ neighbors, i.e., $M^{i}_{k}\hskip-0.85358pt<\hskip-0.85358ptM$, weights for $m\hskip-0.85358pt>\hskip-0.85358ptM^{i}_{k}$ are set equal to zero. We also considered a learnable attention mechanism based on Query-Key interactions [52], but, in the autonomous driving scenario used for performance evaluation, due to limited training data, the resulting prediction performance was inferior to that of the proposed distance-based rule.
To compute an updated hidden state $\bm{\mathbf{h}}^{i}_{k}$, the feature vectors $f^{i}_{\bm{\mathbf{x}}_{k-1}}$ and $f^{i}_{\bm{\mathbf{s}}_{k-1}}$ are stacked and, together with the previous hidden state $\bm{\mathbf{h}}^{i}_{k-1}$, used as an input to the GRU, i.e.
@@
$\bm{\mathbf{x}}^{i}_{k}=\mathrm{Decoder}(\bm{\mathbf{h}}^{i}_{k}).\vskip-11.38109pt$
@@ IV-C The Neural-Enhanced Prediction Step
The predicted PDF for each legacy PO is obtained via marginalization,
i.e.,
@@
$\hskip 8.53581pt=\hskip-0.85358pt\sum_{r^{i}_{k-1}\in\{0,1\}}\hskip-0.85358pt\int\int\int f(\underline{\bm{\mathbf{x}}}^{i}_{k},\underline{r}^{i}_{k}|\bm{\mathbf{x}}^{i}_{k-1},r^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1})$
$\hskip 8.53581pt\times f(\bm{\mathbf{y}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1}|\bm{\mathbf{z}}_{1:k-1})\mathrm{d}\bm{\mathbf{x}}^{i}_{k-1}\mathrm{d}{\mathbf{s}}^{i}_{k-1}\mathrm{d}\bm{\mathbf{h}}^{i}_{k-1}.$ (22)
For $\underline{r}^{i}_{k}\!\hskip-0.85358pt=\!0$, we have $f(\underline{\bm{\mathbf{x}}}^{i}_{k},0\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1})=p\big(\underline{r}^{i}_{k}\!\hskip-0.85358pt=\hskip-0.85358pt0|\bm{\mathbf{z}}_{1:k-1}\big)$ $f_{\mathrm{d}}\big(\underline{\bm{\mathbf{x}}}^{i}_{k}\big)$, where $p\big(\underline{r}^{i}_{k}\!\hskip-0.85358pt=\hskip-0.85358pt0|\bm{\mathbf{z}}_{1:k-1}\big)$ can be obtained as
$p\big(\underline{r}^{i}_{k}\!\hskip-0.85358pt=\hskip-0.85358pt0|\bm{\mathbf{z}}_{1:k-1}\big)$
$\hskip 2.84526pt=\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt\int\hskip-0.85358pt\hskip-0.85358ptf(\bm{\mathbf{x}}^{i}_{k-1},0|\bm{\mathbf{z}}_{1:k-1})+\!\big(1\!-\hskip-0.85358ptp_{\mathrm{s}}\big)f(\bm{\mathbf{x}}^{i}_{k-1},1|\bm{\mathbf{z}}_{1:k-1})\,\mathrm{d}\bm{\mathbf{x}}^{i}_{k-1}$
$\hskip 2.84526pt=p\big(r^{i}_{k-1}\!\hskip-0.85358pt=\hskip-0.85358pt0|\bm{\mathbf{z}}_{1:k-1}\big)+\!\big(1\!-\hskip-0.85358ptp_{\mathrm{s}}\big)p\big(r^{i}_{k-1}\!\hskip-0.85358pt=\hskip-0.85358pt1|\bm{\mathbf{z}}_{1:k-1}\big).$
Here, we used that fact for $\underline{r}^{i}_{k}\!\hskip-0.85358pt=\hskip-0.85358pt0$, the state transition function in (22) simplifies as $f(\underline{\bm{\mathbf{x}}}^{i}_{k},\underline{r}^{i}_{k}\!\hskip-0.85358pt=\hskip-0.85358pt0|\bm{\mathbf{x}}^{i}_{k-1},r^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1})=f(\underline{\bm{\mathbf{x}}}^{i}_{k},\underline{r}^{i}_{k}\!\hskip-0.85358pt=\hskip-0.85358pt0|\bm{\mathbf{x}}^{i}_{k-1},r^{i}_{k-1})$ (cf. (6) and (7)).
For $\underline{r}^{i}_{k}\!\hskip-0.85358pt=\!1$, the marginalization in (22)
reads
$f(\underline{\bm{\mathbf{x}}}^{i}_{k},1\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1})$
$\hskip 2.84526pt=p\big(r^{i}_{k-1}\!\hskip-0.85358pt=\hskip-0.85358pt1|\bm{\mathbf{z}}_{1:k-1}\big)\hskip 0.85358pt\hskip 0.85358ptp_{\mathrm{s}}\hskip-0.85358pt\int\hskip-0.85358pt\int\hskip-0.85358pt\int\hskip-0.85358ptf({\underline{\bm{\mathbf{x}}}}^{i}_{k}|{\bm{\mathbf{x}}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1})$
$\hskip 2.84526pt\times f({\bm{\mathbf{x}}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1},r^{i}_{k-1}\!\hskip-0.85358pt=\hskip-0.85358pt1)\hskip 0.85358pt\mathrm{d}\bm{\mathbf{x}}^{i}_{k-1}\mathrm{d}{\mathbf{s}}^{i}_{k-1}\mathrm{d}{\mathbf{h}}^{i}_{k-1}.$
@@
$f(\bm{\mathbf{h}}^{i}_{k}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1},r^{i}_{k-1}\!\hskip-0.85358pt=\hskip-0.85358pt1)$
@@
$\hskip 1.42262pt\times f({\bm{\mathbf{x}}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1},r^{i}_{k-1}\!\hskip-0.85358pt=\hskip-0.85358pt1)\hskip 0.85358pt\mathrm{d}\bm{\mathbf{x}}^{i}_{k-1}\mathrm{d}{\mathbf{s}}^{i}_{k-1}\mathrm{d}{\mathbf{h}}^{i}_{k-1}.$
@@
To implement (LABEL:eq:prediction_1Prop) and (LABEL:eq:prediction_1Proph), we furthermore approximate the marginal posterior PDF $f({\bm{\mathbf{x}}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1})$
as
$f({\bm{\mathbf{x}}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1},r^{i}_{k-1}\!\hskip-0.85358pt=\hskip-0.85358pt1)$
$\approx f({\bm{\mathbf{x}}}^{i}_{k-1}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1},r^{i}_{k-1}\!\hskip-0.85358pt=\hskip-0.85358pt1)\hskip 0.85358ptf(\bm{\mathbf{h}}^{i}_{k-1}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1},r^{i}_{k-1}\!\hskip-0.85358pt=\hskip-0.85358pt1)$
$\hskip 79.66771pt\times\prod^{M}_{m=1}f(\bm{\mathbf{s}}^{(i,m)}_{k-1}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1},r^{i}_{k-1}\!\hskip-0.85358pt=\hskip-0.85358pt1)$
$\approx\mathcal{N}\big(\bm{\mathbf{\mu}}_{{\bm{\mathbf{x}}}^{i}_{k-1}}\hskip-0.85358pt,\bm{\mathbf{C}}_{{\bm{\mathbf{x}}}^{i}_{k-1}}\big)\hskip 0.85358pt\delta\big(\bm{\mathbf{h}}^{i}_{k-1}-\hat{\bm{\mathbf{h}}}^{i}_{k-1}\big)$
$\hskip 79.66771pt\times\prod^{M}_{m=1}\mathcal{N}\big(\bm{\mathbf{\mu}}_{{\bm{\mathbf{s}}}^{(i,m)}_{k-1}}\hskip-0.85358pt,\bm{\mathbf{C}}_{{\bm{\mathbf{s}}}^{(i,m)}_{k-1}}\big)$, (25)
where $\delta(\cdot)$ denotes the Dirac delta function. Computation of the point estimate $\hat{\bm{\mathbf{h}}}^{i}_{k-1}$ and the Gaussian parameters $\bm{\mathbf{\mu}}_{{\bm{\mathbf{x}}}^{i}_{k-1}}$, $\bm{\mathbf{C}}_{{\bm{\mathbf{x}}}^{i}_{k-1}}\hskip-0.85358pt$, and $\bm{\mathbf{\mu}}_{{\bm{\mathbf{s}}}^{(i,m)}_{k-1}}$, $\bm{\mathbf{C}}_{{\bm{\mathbf{s}}}^{(i,m)}_{k-1}}$, $m\in\{1,\dots,M\}$ will be discussed in the last paragraph of this section. Based on the approximation in (25), we can implement (LABEL:eq:prediction_1Prop) and (LABEL:eq:prediction_1Proph) by means of SPs. An SP-based implementation is motivated by the nonlinear nature of the motion model $g(\bm{\mathbf{x}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1})$ and the fact that an implementation based on particles would be computationally infeasible. Note that the fact that the joint PDF $f({\bm{\mathbf{x}}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1})$ fully factorizes as in (25), is not required for an SP-based implementation. This factorization naturally arises from the computation of marginal posterior PDFs by means of BP [38].
For SP-based implementation of the high-dimensional integration in (LABEL:eq:prediction_1Prop) and (LABEL:eq:prediction_1Proph), we first define a stacked mean vector $\bm{\mathbf{\mu}}^{i}_{k-1}\hskip-0.85358pt\in\hskip-0.85358pt\mathbb{R}^{L(M+1)}$ and covariance matrix $\bm{\mathbf{C}}^{i}_{k-1}\hskip-0.85358pt\in\hskip-0.85358pt\mathbb{R}^{L(M+1)\times L(M+1)}$ corresponding to the product $\mathcal{N}\big(\bm{\mathbf{\mu}}_{{\bm{\mathbf{x}}}^{i}_{k-1}}\hskip-0.85358pt,\bm{\mathbf{C}}_{{\bm{\mathbf{x}}}^{i}_{k-1}}\big)\prod^{M}_{m=1}\mathcal{N}\big(\bm{\mathbf{\mu}}_{{\bm{\mathbf{s}}}^{(i,m)}_{k-1}}\hskip-0.85358pt,\bm{\mathbf{C}}_{{\bm{\mathbf{s}}}^{(i,m)}_{k-1}}\big)$ [36], i.e.
@@
$\hskip-2.84526pt\bm{\mathbf{C}}^{i}_{k-1}$ $\hskip 0.85358pt\triangleq\,\mathrm{bdiag}\Big\{\bm{\mathbf{C}}_{{\bm{\mathbf{x}}}^{i}_{k-1}},\bm{\mathbf{C}}_{{\bm{\mathbf{s}}}^{(i,1)}_{k-1}},\bm{\mathbf{C}}_{{\bm{\mathbf{s}}}^{(i,2)}_{k-1}},\ldots,\bm{\mathbf{C}}_{{\bm{\mathbf{s}}}^{(i,M)}_{k-1}}\Big\}$.
@@
From $\bm{\mathbf{\mu}}^{i}_{k-1}$ and $\bm{\mathbf{C}}^{i}_{k-1}$, $P\hskip-0.85358pt=\hskip-0.85358pt1\hskip-0.85358pt+\hskip-0.85358pt2L(M+1)$ weighted SPs $\big\{\big(\omega^{(i,p)}_{k}\hskip-0.85358pt,\bm{\mathbf{x}}^{(i,p)}_{k-1},{\mathbf{s}}^{(i,p)}_{k-1}\big)\big\}^{P}_{p=1}$ can be computed [20, 36]. SPs are next propagated through the augmented
function, i.e, $\big[\bm{\mathbf{x}}^{(i,p)}_{k}\hskip-0.85358pt\hskip-0.85358pt,{\mathbf{h}}^{(i,p)}_{k}\big]=g\big(\bm{\mathbf{x}}^{(i,p)}_{k-1},{\mathbf{s}}^{(i,p)}_{k-1},\hat{\bm{\mathbf{h}}}^{i}_{k-1}\big)$, $p\hskip-0.85358pt\in\hskip-0.85358pt\{1,\dots,P\}$
. In this way
, we obtain SPs $\big\{\big(\omega^{(i,p)}_{k}\hskip-0.85358pt\hskip-0.85358pt,\bm{\mathbf{x}}^{(i,p)}_{k}\big)\big\}^{P}_{p=1}$ and $\big\{\big(\omega^{(i,p)}_{k}\hskip-0.85358pt\hskip-0.85358pt,{\mathbf{h}}^{(i,p)}_{k}\big)\big\}^{P}_{p=1}$.
Based on these SPs, a Gaussian approximation of $f(\underline{\bm{\mathbf{x}}}^{i}_{k},1|\bm{\mathbf{z}}_{1:k-1})$, is finally obtained as
$f(\underline{\bm{\mathbf{x}}}^{i}_{k},1|\bm{\mathbf{z}}_{1:k-1})\approx p\big(r^{i}_{k-1}\!\hskip-0.85358pt=\hskip-0.85358pt1|\bm{\mathbf{z}}_{1:k-1}\big)\hskip 0.85358ptp_{\mathrm{s}}\hskip 0.85358pt\hskip 0.85358pt\mathcal{N}\big(\bm{\mathbf{\mu}}_{{\underline{\bm{\mathbf{x}}}}^{i}_{k}}\hskip-0.85358pt,\bm{\mathbf{C}}_{\underline{{\bm{\mathbf{x}}}}^{i}_{k}}\big).$
Here, mean $\bm{\mathbf{\mu}}_{{\underline{\bm{\mathbf{x}}}}^{i}_{k}}$ and the covariance matrix $\bm{\mathbf{C}}_{\underline{{\bm{\mathbf{x}}}}^{i}_{k}}$ are computed
following
$\bm{\mathbf{\mu}}_{{\underline{\bm{\mathbf{x}}}}^{i}_{k}}\hskip-0.85358pt=$ $~\sum_{p=1}^{P}\hskip-0.85358pt\omega^{(i,p)}_{k}\bm{\mathbf{x}}^{(i,p)}_{k}$
$\bm{\mathbf{C}}_{\underline{{\bm{\mathbf{x}}}}^{i}_{k}}\hskip-0.85358pt=$ $~{\bm{\mathbf{Q}}}\hskip-0.85358pt+\hskip-0.85358pt\sum_{p=1}^{P}\hskip-0.85358pt\omega^{(i,p)}_{k}\hskip-0.85358pt(\bm{\mathbf{x}}^{(i,p)}_{k}\!-\hskip-0.85358pt\bm{\mathbf{\mu}}_{{\underline{\bm{\mathbf{x}}}}^{i}_{k}})(\bm{\mathbf{x}}^{(i,p)}_{k}\!-\hskip-0.85358pt\bm{\mathbf{\mu}}_{{\underline{\bm{\mathbf{x}}}}^{i}_{k}})^{\mathrm{T}}.$
Finally, a point estimate of the hidden state can be obtained as
$\hat{\bm{\mathbf{h}}}^{i}_{k}=$ $~\sum^{P}_{p=1}\omega^{(i,p)}_{k}\hskip 0.85358pt\hskip 0.85358pt\bm{\mathbf{h}}^{(i,p)}_{k}$.